OpenStudy (angela210793):

I need some help....@Amistre,Satellite,Watchmath and whoever who can give a solution to my problem... A(2;5) B(5;-1) C(x;y) A,B and C are in the same line, 2|AC|=5|BC| find x+y=?

OpenStudy (shadowfiend):

Hm. Is there an illustration that goes along with this?

OpenStudy (shadowfiend):

Oh never mind. I didn't notice C(x;y)

OpenStudy (angela210793):

kk :)

OpenStudy (shadowfiend):

Ok. And just to double-check, |AC| is presumably the length of the A-C line segment?

OpenStudy (angela210793):

we r supposed to work with vectors in here

OpenStudy (shadowfiend):

Right, so the magnitude of the vector. Sounds good.

OpenStudy (angela210793):

And B is between A and C

OpenStudy (shadowfiend):

Ok, so really quick we can see that: \[\begin{align}|AB| = \sqrt{(5 - 2)^2 + (-1 - 5)^2}\end{align}\] We can see conclude that: \[\begin{align}|AC| &= \sqrt{(x - 2)^2 + (y - 5)^2}\\ |BC| &= \sqrt{(x - 5)^2 + (y - -1)^2}\end{align}\]

OpenStudy (shadowfiend):

So if 2|AC| = 5|BC| then: \[\begin{align} 2\sqrt{(x-2)^2 + (y-5)^2} &= 5\sqrt{(x - 5)^2 + (y + 1)^2}\\ \left(2\sqrt{(x-2)^2 + (y-5)^2}\right)^2 &= \left(5\sqrt{(x - 5)^2 + (y + 1)^2}\right)^2 \end{align}\] We can square both sides to make those ugly square roots go away: \[4(x - 2)^2 + 4(y - 5)^2 = 25(x - 5)^2 + 25(y + 1)^2\]

OpenStudy (angela210793):

won't u have -x^2-66x-12y-85?

OpenStudy (angela210793):

or smth like this

OpenStudy (angela210793):

oops w8 I've made a mistake :(

OpenStudy (shadowfiend):

Then we do the annoying multiplication: \[\begin{align} 4(x^2 - 4x + 4) + 4(y^2 - 10y + 25) &= 25(x^2 - 10x + 25) + 25(y^2 + 2y + 1)\\ 4x^2 - 16x + 16 + 4y^2 - 40y + 100 &= 25x^2 - 150x + 625 + 25y^2 + 50y + 25\\ 4x^2 - 16x + 4y^2 - 40y + 116 &= 25x^2 - 150x + 25y^2 + 50y + 650\\ -534 &= 21x^2 - 134x + 21y^2 + 10y\\ \end{align}\] Now the question is how to get x + y from this.

OpenStudy (shadowfiend):

pet22, post your question on the right and wait your turn. Please don't hijack other questions.

OpenStudy (anonymous):

ok

OpenStudy (shadowfiend):

If you repost it I'll get to it in a second.

OpenStudy (shadowfiend):

I wonder if I took the wrong approach here :)

OpenStudy (angela210793):

i got 566 not 534.... anyway i don't think i know how to solve this O.o

OpenStudy (angela210793):

If it may help my teacher has expressed AC as AB+BC so (AB+BC)/BC=5/2 AB/BC=5/2-1=3/2 BC=2AB/3 I got it.....

OpenStudy (angela210793):

Thanks :):):)

OpenStudy (shadowfiend):

Aha. Nice! Rock on :)